# Identifying the magnetic genes in fully- and partially-ordered V2Al ( Cr, Mn, Fe, Co, Ni) Heusler alloys

### v1

@@ Abstract
Multicomponent Heusler alloys exhibit various magnetic properties arising from their diverse atomic compositions and crystal structures. Identifying the general physical principles that govern these behaviors is essential for advancing their potential in spintronic applications. In this work, we combine density functional theory with atomistic Monte Carlo simulations to investigate the magnetic ground states, finite-temperature magnetic transitions, and electronic structures of fully-ordered -, -type, and partially-ordered VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys. We introduce the concept of magnetic genes, defined as V--V triangular motifs connected by the nearest-neighbor (NN) exchange interactions . Within this framework, the magnetic ground states and transition temperatures across the VAl family can be consistently understood. The magnetic order is primarily governed by the NN interactions in the triangular genes, while the transition temperatures are additionally influenced by couplings. Furthermore, the magnetic genes are still proven to be effective in our calculations on partially-ordered V2Al alloys from to -type structures. Our results suggest that the concept of magnetic genes provides a unifying principle for understanding magnetic ordering in V-based Heusler alloys and could serve as a powerful guide for exploring magnetism and designing advanced spintronic materials in a broader class of Heusler systems.
@@ I INTRODUCTION
Although the magnetic properties of full-Heusler alloys have been extensively investigated experimentally, the underlying mechanisms remain elusive due to their complex atomic compositions, multi-sublattice character, and the presence of chemical disorder. Using first-principles approaches, previous works have systematically evaluated the magnetic exchange couplings for several Heusler alloy families employing frozen-magnon and linear-response techniques [28, 29]. For vanadium-based antiferromagnets, existing studies have primarily concentrated on the ordered -type structure of VAl ( Nb, Ta) [30]. Follow-up investigations extended this work by comparing the phase with the inverse phase across a broader range of compounds. In particular, Kuroda et al. reported that -type V2NbAl and V2TaAl favor antiferromagnetic coupling, whereas in VSi ( Ti, Zr, Hf), the structure often competes with or destabilizes the phase [30]. In addition, large-scale first-principles screenings of the inverse-Heusler chemical space, where Sc, Ti, V, Cr, Mn; Al, Si, As; and spans from Ti to Zn, have identified the necessary conditions for the emergence of semimetallic states and established three Slater–Pauling rules associated with – hybridization [31]. Despite these advances, the systematic experimental synthesis and theoretical calculations of V-based VAl compounds with the complete set of Cr, Mn, Fe, Co, and Ni are still lacking. A comprehensive microscopic understanding of the mechanisms that govern magnetism in both fully- and partially-ordered V-based full-Heusler alloys is therefore essential.
@@
In this work, we investigate the ground-state magnetic orders and electronic structures of VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys using density functional theory. For the fully-ordered -type VAl, V2CrAl and V2MnAl exhibit ferrimagnetic order, whereas the remaining compounds adopt ferromagnetic order. The situation becomes more intricate in the inverse -type VAl. While V2CrAl displays a nearly vanishing net magnetic moment with ferromagnetic order between the two different Wyckoff sites of V atoms, the other four alloys exhibit ferrimagnetic order with antiparallel spin alignment between the two V sublattices. Despite this diversity, the magnetic orders across VAl compounds can be consistently interpreted in terms of the V--V triangular pathway, governed by the NN magnetic exchange coupling . Furthermore, finite-temperature magnetic transitions were systematically investigated through atomistic Monte Carlo simulations, which reveal that the sequence of Curie temperatures among different compositions and lattice types is primarily determined by , with additional contributions from exchange interactions. Importantly, we find that the V--V triangular pathway framework remains valid even in partially-ordered systems from to -type structures. Building on this insight, we propose the concept of “magnetic genes”, defined as the V--V triangular motif connected by NN exchange interactions . We recognize that the magnetic genes could provide a promising concept for exploring the magnetic properties in more Heusler alloys.
@@ II METHODS
To perform the structural relaxation calculations, we use the VASP code [32, 33], which implements the projector-augmented wave (PAW) method [34] and is based on the GGA-PBE generalized gradient approximation [35]. A dense -mesh of 10×10×10 is employed for the structural relaxation calcuations. The plane-wave cutoff was chosen to be 425 eV. The convergence criteria for the total energy and ionic forces were set to and , respectively.
Additionally, we performed spin-polarized density-functional [36, 37] calculations using the Atomic-orbital Based Ab-initio Computation at USTC (ABACUS) package [38, 39], an open-source first-principles platform with advanced ground- and excited-state modules, and interfaces to Wannier90. We employed the SG15 optimized norm-conserving Vanderbilt (ONCV) pseudopotentials [40] and the same exchange correlation [35]. The Kohn-Sham wave functions were expanded using a linear combination of atomic orbitals corresponding to the SG15 ONCV (Version 2.0 standard orbitals) [39]. After convergence tests, the grid cutoff for numerical integrations was set to 100 Ry.
@@
The simulated system of V2Al is a cubic box of edge length , generated by replicating the relaxed face-centered unit cell and containing about spins, which was evolved with a stochastic spin integrator. Temperatures from to in increments were sampled using equilibration steps. The Curie temperature was determined from the inflection point of the magnetization .
@@ III RESULTS AND DISCUSSION
III.1 Magnetic orders of fully-ordered - and -V2Al
The magnetic ordering in Heusler alloys is strongly influenced by both their atomic composition and crystallographic structure. In this work, we focus on two prototypical lattice types, namely the and structures. As established in previous studies, these two structural variants provide a well-defined platform for examining the role of chemical disorder without altering the overall structural topology [7, 14, 28]. As illustrated in Fig. 1(a), the -type full-Heusler structure (space group , No.225) with the generic formula VAl consists of V atoms occupying the 8 Wyckoff positions, while Al and atoms are located at the 4 and 4 sites, respectively [7]. By contrast, in the -type inverse-Heusler structure (space group , No.216), shown in Fig. 1(b), the atoms occupy the 4 positions that arise from the symmetry breaking of the 8 sites in the -type structure, while the V atoms are shifted to the 4 sites. From a broader perspective, the redistribution of atomic occupations between the 4 and 4 sites, driven by the evolution from to -type structures, highlights the importance of chemical disorder in determining the magnetic properties of these compounds. A detailed analysis of disorder effects will be presented in Sec. III.4.
@@ III.1 Magnetic orders of fully-ordered - and -V2Al
We begin by examining the effect of atomic composition on the magnetic order in -type VAl ( Cr, Mn, Fe, Co, Ni), where the crystal structure is relatively simple, containing only three inequivalent Wyckoff sites compared to the more complex -type structure. As summarized in Table 1, compounds with Fe, Co, and Ni exhibit ferromagnetic ordering, whereas those with Cr and Mn display ferrimagnetic ordering characterized by antiparallel spin alignment between the and V sublattices. In particular, our calculations show that V2MnAl adopts a ferrimagnetic ground state, with magnetic moments of per V atom and per Mn atom. By contrast, V2FeAl exhibits ferromagnetic ordering with a total magnetic moment of per formula unit, in agreement with previous reports [43]. Notably, the total magnetic moments of the Fe to Ni compounds do not vary monotonically across the series. While the magnetic moments on the V sublattice increase progressively from Fe to Ni, those on the site decrease rapidly, leading to non-monotonic behavior in the total moment. These complex magnetic orders and the non-monotonic variation in total magnetization are primarily attributed to differences in the underlying electronic structures arising from the atomic composition. A detailed discussion of the interplay between electronic structure and magnetic order will be presented in Sec. III.5.
@@
III.2 Microscopic magnetic mechanism in - and -V2Al
@@
To gain deeper insight into the origin of magnetism in these systems, we investigate the magnetic exchange couplings () across different atomic compositions and lattice types. As discussed above, in the relatively simple -type VAl compounds, only those with Cr and Mn exhibit ferrimagnetic order, characterized by anti-parallel spin alignment between the and V atoms. Correspondingly, as shown in Figs. 2(a) and 2(b), the NN exchange coupling () between Mn (Cr) atoms at the 4 site and V atoms at the 8 site is antiferromagnetic type, which has the largest value among all the other magnetic exchange couplings in these materials. In contrast, for Fe, Co, and Ni, the NN coupling between the and V atoms is ferromagnetic, as illustrated in Fig. 2(c). Meanwhile, the exchange interactions between - atoms are significantly weaker and do not play a decisive role in stabilizing the magnetic order. The V-V interactions, however, exhibit more intricate behavior. In V2MnAl, both the NN and next-nearest-neighbor (NNN) V-V couplings, indicated by the gray line in Fig. 2(a), are ferromagnetic. However, the NN V-V coupling in V2CrAl, colored by the gray line, acquires a finite antiferromagnetic contribution, as shown in Fig. 2(b). Taken together, these results indicate that antiferromagnetic V-Mn(Cr)-V exchange pathways form triangular motifs that ultimately stabilize ferromagnetic ordering among the V sublattice. Overall, our exchange-coupling analysis provides a microscopic explanation for the emergence of complex magnetic orders in -type VAl compounds.
@@
Here, we investigate the physical mechanisms underlying the magnetic ordering in the more complex -type VAl compounds. In these type compounds, the magnetic moments of V1 (4 site) and (4 site) are aligned antiparallel, indicating dominant antiferromagnetic exchange coupling. As shown in Figs. 2(d) and 2(e), the NN exchange couplings (highlighted in blue) are strongly antiferromagnetic ( Cr, Mn; couplings for Fe, Co, Ni are omitted for clarity) and represent the primary factor stabilizing ferrimagnetic ordering between the and V1 sublattices. In contrast, the NN exchange coupling between and V2 (4 site), as shown in purple lines in Figs. 2(e) and 2(f), evolves rapidly from antiferromagnetic to ferromagnetic as changes from Cr to Ni. This change of is consistent with the corresponding magnetic ordering from ferrimagnetic for Cr to ferromagnetic for Mn, Fe, Co, and Ni. Notably, although the NN V2-Mn interaction is weakly antiferromagnetic type, the NNN V2-Mn coupling is strongly ferromagnetic, suggesting that V2MnAl lies near a compositional boundary where the magnetic ordering transitions from ferrimagnetic to ferromagnetic with increasing atomic number.
@@
From our analysis of magnetic orders and exchange couplings, a comprehensive understanding of the overall magnetic behavior in VAl across different atomic compositions and lattice types can be established. In -type VAl, the diverse magnetic orders can be rationalized in terms of triangular exchange pathways formed by V--V atoms, as illustrated in the inset of Fig. 2(a). In -type VAl, the underlying mechanism is similar, but an additional type of magnetic exchange coupling is present. Our results indicate that the V1--V2 triangular pathway remains a key determinant of magnetic order in -type compounds, as shown in the inset of Fig. 2(d). Changes in the type of the coupling substantially influence the magnetic ordering in these materials. Moreover, although the direct NN exchange coupling between - atoms ( Mn, Fe, Co, Ni) is antiferromagnetic type in -VAl, the triangular -V1- pathway ultimately stabilizes ferromagnetic ordering due to the intervening antiferromagnetic X-V1 couplings, as illustrated in the inset of Fig. 2(d). In summary, despite variations in atomic composition and lattice type, the site plays a central role in controlling the magnetic properties of V-based VAl compounds through triangular pathways formed by NN V and atoms. We therefore identify these motif as a candidate “magnetic genes”.
III.3 Magnetic transition temperature of - and -V2Al

### v2

# Local magnetic structure in fully and partially ordered V2Al Heusler alloys (=Cr, Mn, Fe, Co, Ni)

@@ Abstract
Multicomponent Heusler alloys exhibit various magnetic properties arising from their diverse atomic compositions and crystal structures. Identifying the general physical principles that govern these behaviors is essential for advancing their potential in spintronic applications. In this work, we combine density functional theory with atomistic Monte Carlo simulations to investigate the magnetic ground states, finite-temperature magnetic transitions, and electronic structures of fully-ordered -, -type, and partially-ordered VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys. We propose the concept of magnetic motifs, defined as V--V triangular pathway connected by the nearest-neighbor (NN) exchange interactions . Within this framework, the magnetic ground states and transition temperatures across the VAl family can be consistently understood. The magnetic order is primarily governed by the NN interactions in the triangular motifs, while the transition temperatures are additionally influenced by couplings. Furthermore, the magnetic motifs are still proven to be effective in our calculations on partially-ordered V2Al alloys from to -type structures. Our results suggest that the concept of magnetic motifs provides a unifying principle for understanding magnetic ordering in V-based Heusler alloys and could serve as a candidate guide for exploring magnetism and designing advanced spintronic materials in a broader class of Heusler systems.
@@ I INTRODUCTION
Although the magnetic properties of full-Heusler alloys have been extensively investigated experimentally, the underlying mechanisms remain elusive due to their complex atomic compositions, multi-sublattice character, and the presence of chemical disorder. Using first-principles approaches, previous works have systematically evaluated the magnetic exchange couplings for several Heusler alloy families employing frozen-magnon and linear-response techniques [9, 3]. For vanadium-based antiferromagnets, existing studies have primarily concentrated on the ordered -type structure of VAl ( Nb, Ta) [26]. Follow-up investigations extended this work by comparing the phase with the inverse phase across a broader range of compounds. In particular, Kuroda et al. reported that -type V2NbAl and V2TaAl favor antiferromagnetic coupling, whereas in VSi ( Ti, Zr, Hf), the structure often competes with or destabilizes the phase [26]. Despite these advances, a synthetically theoretical investigation of magentism of V element in VAl compounds remains limited, especially when is considered as a magnetic element. A comprehensive microscopic understanding of the mechanisms that govern magnetism in both fully- and partially-ordered V-based full-Heusler alloys is therefore essential.
@@
In this work, we investigate the ground-state magnetic orders and electronic structures of VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys using density functional theory. V atoms play a key role in regulating the magnetic moments of the five magnetic transition elements, influencing both the orientation and magnitude of the magnetic moments. For the fully-ordered -type VAl, V2CrAl and V2MnAl exhibit ferrimagnetic order, whereas the remaining compounds adopt ferromagnetic order. The situation becomes more intricate in the inverse -type VAl. While V2CrAl displays a nearly vanishing net magnetic moment with ferromagnetic order between the two V atoms on different Wyckoff sites, the other four alloys exhibit ferrimagnetic order with antiparallel spin alignment between the two V sublattices. Despite this diversity, the magnetic orders across VAl compounds can be consistently interpreted in terms of the V--V triangular pathway, governed by the NN magnetic exchange coupling . Furthermore, finite-temperature magnetic transitions were systematically investigated through atomistic Monte Carlo simulations, which reveal that the sequence of Curie temperatures among different compositions and lattice types is primarily determined by , with additional contributions from exchange interactions. Importantly, we find that the V--V triangular pathway framework remains valid even in partially-ordered systems from to -type structures. Building on this insight, we propose the concept of “magnetic motifs”, defined as the V--V triangular motif connected by NN exchange interactions . We recognize that the magnetic motifs could provide a promising concept for exploring the magnetic properties in more Heusler alloys.
@@ II METHODS
All calculations were performed based on spin-polarized density functional theory [15, 21]. To get the equilibrium-state structures of Heusler alloys, we used the VASP code to relax the crystal structures [22, 23], which implements the projector-augmented wave (PAW) method [2], and the generalized gradient approximation in the form of Perdew-Burke-Ernzerhof (GGA-PBE) functional was employed to represent the exchange-correlation potential with a dense -mesh of 10×10×10. The plane-wave cutoff was chosen to be 425 eV. The convergence criteria for the total energy and ionic forces were set to and , respectively.For each composition of VAl, we first obtained the relaxed lattice constants of the fully-ordered structure. Based on these optimized configurations, partially-ordered structures were constructed by exchanging V and atoms, as illustrated in Fig. 1(c). Each partially ordered model was subsequently subjected to an independent lattice relaxation. In this study, all results are obtained from the standard DFT calculations without explicitly accounting for strong electronic correlations. Previous DFT investigations have shown that exchange interactions and Curie temperatures in many Heusler alloys can be accurately reproduced without the inclusion of a Hubbard term, yielding good agreement with experimental measurements [24, 40, 35]. For comparison, the results of DFT+ calculations are provided in the Supplementary Information [1] (see also references [29, 45, 55] therein). While the inclusion of the Coulomb interaction does not change the main physical trends reported here, the DFT+ approach generally overestimates the magnitude of the exchange interactions, leading to unrealistically high Curie temperatures.
We calculated the magnetic orders and electronic structures by using the Atomic-orbital Based Ab-initio Computation at USTC (ABACUS) package [27, 56], an open-source first-principles platform with advanced ground- and excited-state modules, and interfaces to Wannier90. We employed the SG15 optimized norm-conserving Vanderbilt (ONCV) pseudopotentials [42] and the same exchange correlation functional GGA-PBE [36]. The Kohn–Sham electron wave functions were expanded in terms of a linear combination of atomic orbitals (LCAO) associated with the SG15 ONCV pseudopotentials (2.0) [56]. After convergence tests, the grid cutoff for numerical integrations was set to 100 Ry.
@@
In Monte Carlo simulations, we examined the convergence with respect to both the supercell size and -point sampling. We found that the magnetic transition temperature exhibited negligible variation for supercell sizes ranging from 10 to 20 nm. Similarly, for -point grids ranging from to , the results were essentially unchanged for grids finer than . Based on these convergence tests, we selected the Monte Carlo supercell of 15 nm and the -point mesh of for all subsequent calculations. The simulated system of V2Al is a cubic box of edge length 15 nm, generated by replicating the relaxed face-centered unit cell and containing about spins, which was evolved with a stochastic spin integrator. Temperatures from to in increments were sampled using equilibration steps. The Curie temperature was determined from the inflection point of the magnetization .
@@ III RESULTS AND DISCUSSION
III.1 Structural configuration and thermodynamic stability analysis of V2Al
The magnetic ordering in Heusler alloys is strongly influenced by both their atomic composition and crystallographic structure. In this work, we focus on two prototypical lattice types, namely the and structures. As established in previous studies, these two structural variants provide a well-defined platform for examining the role of chemical disorder without altering the overall structural topology [11, 37, 9]. As illustrated in Fig. 1(a), the -type full-Heusler structure (space group , No.225) with the generic formula VAl consists of V atoms occupying the 8 Wyckoff positions, while Al and atoms are located at the 4 and 4 sites, respectively [11]. By contrast, in the -type inverse-Heusler structure (space group , No.216), shown in Fig. 1(b), the atoms occupy the 4 positions that arise from the symmetry breaking of the 8 sites in the -type structure, while the V atoms are shifted to the 4 sites. From a broader perspective, the redistribution of atomic occupations between the 4 and 4 sites, driven by the evolution from to -type structures, highlights the importance of chemical disorder in determining the magnetic properties of these compounds. The details of the relaxed lattice parameters for all configurations are summarized in Supplementary Information. A detailed analysis of disorder effects will be presented in Sec. III.5.
In the VAl ( Cr, Mn, Fe, Co, Ni) Heusler alloys, differences in electronegativity among the constituent atoms play a critical role in determining chemical bonding and, consequently, the structural stability of the compounds. The electronegativity difference between vanadium and the -site elements increases from 0.03 to 0.29 Pauli units as varies from Cr to Ni, enhancing the strength of the V- chemical bonds and contributing to improved thermodynamic stability across the series. Similarly, the electronegativity difference between Al and the -site atoms rises from 0.05 to 0.31 Pauli units along the same series, providing additional stabilization through -Al chemical bonding. In contrast, the electronegativity difference between V and Al is relatively small (0.02 Pauli units), lower than that of both the V- and -Al pairs. Nevertheless, this difference still introduces supplementary bonding that help stabilize the overall Heusler structure.
To assess the stability of the fully ordered and partially ordered V2Al compounds, we calculated both the formation energies and total energies of the possible structural configurations. As shown in Fig. 2(a), all systems, except for the -type V2NiAl, exhibit negative formation energies, indicating their thermodynamic stability. These results are consistent with the trends reported in the Open Quantum Materials Database (OQMD) [39, 20]. The calculated formation energies further reveal that V2MnAl, V2FeAl, and V2CoAl preferentially adopt the -type structure, whereas V2CrAl is most stable in the configuration. For V2NiAl, a partially disordered atomic arrangement is energetically favored. These conclusions align well with previous theoretical and experimental studies on V-based Heusler alloys [32, 55, 47]. The corresponding total energy results for both the fully-ordered structures and the partially-ordered structural pathways are presented in Fig. 2(b). Consistent with the formation energy trends, the -type ordering is the lowest-energy configuration for both V2MnAl and V2FeAl. Moreover, the total energies of the partially-ordered configurations lie between those of the parent - and -type phases. These findings underscore the importance of considering phase stability as a key criterion in the practical design and optimization of high-performance Heusler-based spintronic devices.
Here, we note that the partially-ordered Heusler alloys are constructed by progressively exchanging V atoms at the 4 sites with atoms at the 4 sites within a 16-atom cubic cell. This procedure generates a series of partially-ordered “intermediate” structures. However, we find that different disorder configurations for a given composition do not lead to much variation in the exchange interactions. Therefore, in this study, we present the results for the configuration with the lowest total energy.
@@ III.1 Structural configuration and thermodynamic stability analysis of V2Al
III.2 Magnetic orders of fully-ordered - and -V2Al
We begin by examining the effect of atomic composition on the magnetic order in -type VAl ( Cr, Mn, Fe, Co, Ni), where the crystal structure is relatively simple, containing only three inequivalent Wyckoff sites compared to the more complex -type structure. As summarized in Table 1, compounds with Fe, Co and Ni exhibit ferromagnetic ordering, whereas those with Cr and Mn display ferrimagnetic ordering characterized by antiparallel spin alignment between the and V sublattices. In particular, our calculations show that V2MnAl adopts a ferrimagnetic ground state, with magnetic moments of per V atom and per Mn atom. By contrast, V2FeAl exhibits ferromagnetic ordering with a total magnetic moment of per formula unit, in agreement with previous reports [10]. Notably, the total magnetic moments of the Fe to Ni compounds do not vary monotonically across the series. While the magnetic moments on the V sublattice increase progressively from Fe to Ni, those on the site decrease rapidly, leading to non-monotonic behavior in the total moment. These complex magnetic orders and the non-monotonic variation in total magnetization are primarily attributed to differences in the underlying electronic structures arising from the atomic composition. A detailed discussion of the interplay between electronic structure and magnetic order will be presented in Sec. III.6.
@@ III.2 Magnetic orders of fully-ordered - and -V2Al
III.3 Microscopic magnetic mechanism in - and -V2Al
@@
To gain deeper insight into the origin of magnetism in these systems, we investigate the magnetic exchange couplings () across different atomic compositions and lattice types. As discussed above, in the relatively simple -type VAl compounds, only those with Cr and Mn exhibit ferrimagnetic order, characterized by anti-parallel spin alignment between the and V atoms. Correspondingly, as shown in Figs. 3(a) and 3(b), the NN exchange coupling () between Mn (Cr) atoms at the 4 site and V atoms at the 8 site is antiferromagnetic type, which has the largest value among all the other magnetic exchange couplings in these materials. In contrast, for Fe, Co and Ni, the NN coupling between the and V atoms is ferromagnetic, as illustrated in Fig. 3(c). Meanwhile, the exchange interactions between - atoms are significantly weaker and do not play a decisive role in stabilizing the magnetic order. The V-V interactions, however, exhibit more intricate behavior. In V2MnAl, both the NN and next-nearest-neighbor (NNN) V-V couplings, indicated by the gray line in Fig. 3(a), are ferromagnetic. However, the NN V-V coupling in V2CrAl, colored by the gray line, acquires a finite antiferromagnetic contribution, as shown in Fig. 3(b). Taken together, these results indicate that antiferromagnetic V-Mn(Cr)-V exchange pathways form triangular motifs that ultimately stabilize ferromagnetic ordering among the V sublattice. Overall, our exchange-coupling analysis provides a microscopic explanation for the emergence of complex magnetic orders in -type VAl compounds.
@@
Here, we investigate the physical mechanisms underlying the magnetic ordering in the more complex -type VAl compounds. In these type compounds, the magnetic moments of V1 (4 site) and (4 site) are aligned antiparallel, indicating dominant antiferromagnetic exchange coupling. As shown in Figs. 3(d) and 3(e), the NN exchange couplings (highlighted in blue) are strongly antiferromagnetic ( Cr, Mn; couplings for Fe, Co, Ni are shown in Supplementary Information) and represent the primary factor stabilizing ferrimagnetic ordering between the and V1 sublattices. In contrast, the NN exchange coupling between and V2 (4 site), as shown in purple lines in Figs. 3(e) and 3(f), evolves rapidly from antiferromagnetic to ferromagnetic as changes from Cr to Ni. This change of is consistent with the corresponding magnetic ordering from ferrimagnetic for Cr to ferromagnetic for Mn, Fe, Co, and Ni. Notably, although the NN V2-Mn interaction is weakly antiferromagnetic type, the NNN V2-Mn coupling is strongly ferromagnetic, suggesting that V2MnAl lies near a compositional boundary where the magnetic ordering transitions from ferrimagnetic to ferromagnetic with increasing atomic number.
@@
From our analysis of magnetic orders and exchange couplings, a comprehensive understanding of the overall magnetic behavior in VAl across different atomic compositions and lattice types can be established. In -type VAl, the diverse magnetic orders can be rationalized in terms of triangular exchange pathways formed by V--V atoms, as illustrated in the inset of Fig. 3(a). In -type VAl, the underlying mechanism is similar, but an additional type of magnetic exchange coupling is present. Our results indicate that the V1--V2 triangular pathway remains a key determinant of magnetic order in -type compounds, as shown in the inset of Fig. 3(d). Changes in the type of the coupling substantially influence the magnetic ordering in these materials. Moreover, although the direct NN exchange coupling between - atoms ( Mn, Fe, Co, Ni) is antiferromagnetic type in -VAl, the triangular -V1- pathway ultimately stabilizes ferromagnetic ordering due to the intervening antiferromagnetic X-V1 couplings, as illustrated in the inset of Fig. 3(d). In summary, despite variations in atomic composition and lattice type, the site plays a central role in controlling the magnetic properties of V-based VAl compounds through triangular pathways formed by NN V and atoms. We therefore identify these pathway as a candidate “magnetic motifs”.
III.4 Magnetic transition temperature of - and -V2Al